# Latent Concept-based Explanation of NLP Models

### v2

@@ Appendix A Task-specific Prompts used with PlausiFyer
Appendix B Datasets
@@
Appendix C Finetuning Performance
We tuned several transformers BERT-base-cased, RoBERTa and XLM-RoBERTa. We used standard splits for training, development and test data that we used to carry out our analysis. The splits to preprocess the data are available through git repository [^2] https://github.com/nelson-liu/contextual-repr-analysis. See Table 5 and Table 6 for statistics and classifier accuracy. We present the results of Toxicity and MNLI in Appendix H and I.
@@ Appendix C Finetuning Performance
Appendix D Qualitative Evaluation - More Examples
D.1 Example for the Evolution of Concepts
@@
D.2 Adversarial Example of the Sentence in Figure 3(b)
@@
D.3 Correct Predicted Label with Incorrect Gold Label
@@
D.4 Incorrect Prediction in POS tagging Task
@@
Appendix E Module Specific Evaluation
E.1 ConceptDiscoverer - Compositional Concept Examples
@@
E.2 ConceptDiscoverer - Number of Clusters For Each Polarity in the Sentiment Classification Task
@@
E.3 ConceptMapper - Accuracy of ConceptMapper for the Sentiment Classification and POS Tagging task
@@
Appendix F Human Evaluation
F.1 LACOAT Effectiveness
@@
F.2 Comparison with other Methods
For comparison with other methods, we ask four annotators to rank 100 samples where they see the original input, gold label, predicted label, and explanations by three methods: LACOAT, IG and COCKATIEL. LACOAT explanations are shown across three layers (layer 0, 6 and 12), while IG explanations are shown for layer 0 and COCKATIEL for layer 12. The annotators are asked to rank each method from 1 to 3 in terms of usefulness in understanding the reason for the prediction where 1 implies the method was very useful while 3 implies it was not useful. The annotation allows for the annotator to rank multiple methods with the same usefulness rating, e.g. for a particular sample, both LACOAT and COCKATIEL can have the rank 1. This setting is intentional since the output of explanation methods is not directly comparable to each other due to the difference in their design and the targeted form and granularity of explanation. Table 1 presents the results. The results suggested that LACOAT is preferred or equally preferred by all annotators. The average Cohen’s $\kappa$ further shows a "fair agreement" between annotators and the consolidated ranking where consolidated ranking is the average rank across users.
Appendix G Faithfulness Evaluation
We ablated the most salient latent concept for a prediction by subtracting its average representation from the [CLS] representation of layer 12. Random represents the subtraction of a random vector. We report the average results of five random vectors. Accuracy represents the performance of the model on the test set. Prediction change represents the percentage of predictions that altered after manipulation. The results show that manipulating the [CLS] token representation using the LACOAT vector leads to significant drops in performance and changes in predictions across all datasets. In contrast, random vector manipulations have a minimal impact on the model’s performance and predictions. These results suggest that the LACOAT vector plays a crucial role in the model’s decision-making process. Comparing the results of different datasets, MNLI showed a relatively smaller drop in accuracy when manipulating using the salient latent concept vector. We suspect that this is due to the nature of the MNLI task that requires reasoning over multiple sentences and whose information may be present in multiple latent concepts. Nevertheless, the difference in results from original accuracy and random vector confirms our hypothesis of the faithfulness of latent concepts.
Faithfulness Metrics
Dataset Setting Accuracy % Label Flip
Sentiment Original 96.31 -
LACOAT 55.91 43.98
Random 96.09 0.14
Toxicity Original 91.55 -
LACOAT 51.78 46.44
Random 91.93 0.13
MNLI Original 87.69 -
LACOAT 82.08 8.83
Random 88.12 0.55
Table 12: Faithfulness evaluation using the RoBERTa model. Original is the performance of the model without any manipulation, LACOAT is the performance of the model after subtracting the most salient latent concept vector from the [CLS] vector and Random is the average performance of the model across five random vectors when subtracted from the [CLS] vector
Appendix H Toxicity Classification Task
H.1 Experimental Setup

### v3

@@ 2.4 PlausiFyer
3 Experimental Setup
Data
We use Parts-of-Speech (POS) tagging, toxicity classification (Toxicity), sentiment classification (Sentiment), and natural language inference (NLI) tasks for our experiments. POS is a sequence labeling task, while the other tasks are sequence classification tasks. We use the Penn TreeBank dataset Marcus et al. (1993) for POS, the Jigsaw Toxicity dataset cjadams et al. (2017) for toxicity, the ERASER Movie Reviews dataset Pang and Lee (2004) for sentiment, and the MNLI dataset Wang et al. (2019) for the NLI task. Appendix A provides information about each dataset.
Models
We fine-tune 12-layered pre-trained models; BERT-base-cased  Devlin et al. (2019), RoBERTa-base Liu et al. (2019) and XLM-Roberta  Conneau et al. (2020) using the training datasets of the tasks considered. For Llama-2-7b-chat-hf Touvron et al. (2023), we use the base model without finetuning with zero-shot prompting for each task. We use transformers Wolf et al. (2020) with the default settings and hyperparameters. Task-wise performance of the models is provided in Appendix (Tables 6,  713, and 17).
Module-specific hyperparameters
When extracting the activation [^2] We used the NeuroX toolkit Dalvi et al. (2023). and/or attribution of a word, we average the respective value over the word’s subword units Durrani et al. (2019). We optimize the number of clusters $K$ for each dataset as suggested by Dalvi et al. (2022). We use $K=600$ (POS, Toxicity) and $K=400$ (Sentiment, MNLI) for ConceptDiscoverer.
4 Evaluation
We perform a qualitative evaluation, a human evaluation and a module-level evaluation of LACOAT to measure its correctness and efficacy. We find consistent results across all tasks and models. Due to space limitation, we mainly present the results of POS and Sentiment using the BERT and RoBERTa models in the main paper. The full set of results are presented in Apps. FGH.
4.1 Qualitative Evaluation
4.1.1 Evolution of Concepts
LACOAT generates the explanation for each layer with respect to a prediction. The layer-wise explanation shows the evolution of concepts in making the prediction. Figure 2 shows layers 0, 6 and 12’s latent concept of the most attributed input token for RoBERTa fine-tuned on the sentiment task (see Figure 5 in Appendix for more examples). We found that the initial layer latent concepts do not always align with the sentiment of the input instance and may represent a general language concept. For instance, Figure 2(a) shows the concept of comparative and superlative adjectives of both positive and negative sentiments and is not limited to representing the negative sentiment of the most attributed word. In the middle layers, the latent concepts evolved into concepts that align better with the sentiment of the input sentence. For instance, the latent concept of Figure 2(b) shows a mix of adjectives and adverbs of negative sentiment, i.e. aligned with the sentiment of the input sentence. In the sentiment task, the most attributed word in the last layer is [CLS] which resulted in latent concepts consisting of [CLS] representations of the most related sentences to the input. In such cases, we randomly pick five [CLS] instances from the latent concept and show their corresponding sentences in the figure (see Figure 2(c)). We found that the last layer’s latent concepts are best aligned with the input instance and its prediction and are the most informative explanation of the prediction. In the rest of the paper, we focus our analysis on the explanations generated using the last layer only and perform a human evaluation to evaluate their efficacy and correctness.
4.1.2 Analyzing Last Layer Explanations
Figure 3 presents various examples of LACOAT for both POS tagging and sentiment tasks using BERT. The sentence refers to the input sentence, prediction is the model’s output, and true label is the gold label. The explanation is the final output of LACOAT. Cluster X denotes the latent concept aligned with the most salient word representation at the 12th layer, where X is the cluster ID. For sentiment, we randomly select five [CLS] instances from the latent concept and show their corresponding sentences in the figure.
Correct prediction with correct gold label
Figures 3(a) and 3(c) present a case of correct prediction with latent-concept explanation and human-friendly explanation. The former is harder to interpret especially in the case of sentence-level latent concepts as in Figure 3(a) compared to latent concepts consisting of words (Figure 3(c)). However, in both cases, PlausiFyer highlights additional information about the relation between the latent concept and the input sentence. For example, it captures that the adverbs in Figure 3(c) have common semantics of showing degree or frequency. Similarly, it highlights that the reason for the positive sentiment in Figure 3(a) arises from praising different aspects of a film and its actors and actresses.
Wrong prediction with correct gold label
To study how the explanation would change if it is a correct prediction, we employ TextAttack Morris et al. (2020) to create an adversarial example of the sentence in Figure 3(b) that flips its prediction. The new sentence replaces “laughing” with “kidding” which has a similar meaning but flipped the prediction to a correct prediction. Figure 6 in the Appendix. shows the full explanation of the augmented sentence. With the correct prediction, the latent concept changed and the explanation clearly expresses a negative sentiment “… all express negative opinions and criticisms …” compared to the explanation of the wrongly predicted sentence.
Cross model analysis
Figure 4: Comparing explanation of XLMR (top) and RoBERTa (bottom)
4.2 Human Evaluation
LACOAT Effectiveness
We conduct a human evaluation using four annotators across 50 test samples. Specifically, given an explanation (e.g. Figure 3), all annotators are asked to answer five questions (Q1-Q5) that aimed at evaluating the usefulness of LACOAT. [^3] We provide the evaluation questions in Appendix E. Specifically, Q1 evaluates whether LACOAT attributes the correct concept to a given prediction, while Q2 and Q3 measure the efficacy of LACOAT’s output in helping a user understand the prediction. Q4 and Q5 evaluates the output of PlausiFyer. They specifically separate out the cases where the explanation was accurate but irrelevant to the task at hand.
Comparison with other Methods
Despite the numerous proposed explanation methods comparing them due to the difference in granularity, type of explanations, and methodology. We design a human evaluation where evaluators assign a score between 1 to 3 to each of three explanations generated by IG, LACOAT and Cockatiel Jourdan et al. (2023). The annotation setup allows for ranking multiple methods with the same usefulness rating. A total of 400 annotations were collected using four evaluators, each evaluating all test instances. We provide the details of the evaluation setup and the results in Appendix E. The second part of Table 1 shows the percentage of samples for which each annotator ranked LACOAT as the same or better than both IG and Cockatiel. The consolidated ranking is computed by averaging the ranks across users. The average Cohen’s $\kappa$ indicates Fair agreement between each annotator and the consolidated ranking, demonstrating that the LACOAT explanation is more useful for understanding predictions compared to other methods.
Top Labels Correct Samples Incorrect Samples All Samples
Annotation Fleiss $\kappa$
Q1 Yes/No 28 / 0 20 / 2 48 / 2 0.35
Q2 Helps/Neutral/Hinders 27 / 1 / 0 17 / 5 / 0 44 / 6 / 0 0.41
Q3 Helps/Neutral/Hinders 16 / 10 / 2 1 / 19 / 2 17 / 29 / 4 0.61
Q4 Yes/No 17 / 11 5 / 17 22 / 28 0.47
Q5 Yes/No 17 / 11 6 / 16 23 / 27 0.80
Bottom A1 A2 A3 A4 Consolidated Average Cohen’s $\kappa$
LACOAT $\uparrow$ 85% 72% 77% 87% 89% 0.37
4.3 Module Specific Evaluation
POS Sentiment
Layers BERT RoBERTa BERT RoBERTa Llama2
9 (/20) 92.38 86.97 31.94 99.59 70.63
10 (/24) 92.79 89.64 99.57 99.69 75.64
11 (/28) 93.39 89.95 99.71 99.48 71.30
12 (/32) 93.95 90.04 99.25 99.27 71.02
Table 2: Accuracy of PredictionAttributor in mapping a representation to the correct latent concept. The layer number in the parentheses corresponds to the Llama2-chat model.
4.3.1 ConceptDiscoverer
ConceptDiscoverer identifies latent concepts by clustering the representation. We question whether the discovered latent concepts are a true reflection of the properties that a representation possesses. Using ConceptDiscoverer, we form latent concepts of the last layer and automatically annotate them as described above. We found 87%, 83% and 86% of the latent concepts of BERT, RoBERTa and XLMR that perfectly map to a POS tag respectively. We further analyzed other concepts where 90% of the words did not belong to a single tag. We found them to be of compositional nature i.e. a concept consisting of related semantics like a mix of adjectives and proper nouns about countries such as Swedish and Sweden (Appendix:Figure 9). For sentiment, we found 78%, 95%, 94%, and 67% of the latent concepts of BERT, RoBERTa, XLMR, and Llama2-chat to consist of either Positive or Negative sentences. The high number of class-based clusters of RoBERTa and XLMR show that at the last layer, the majority of their latent space is separated based on these two classes (see Table 8 and Table 21 for detailed results).
Layers (BERT) 0 2 5 10 12
POS Top 1 100 100 99.03 92.67 84.19
Top 2 100 100 99.75 97.89 94.15
Top 5 100 100 99.94 99.68 99.05
Sentiment Top 1 100 100 97.19 83.09 68.24
Top 2 100 100 99.63 92.67 83.24
Top 5 100 100 99.94 97.75 94.24
Layers (Llama2-chat) 0 8 16 24 32
Sentiment Top 1 97.55 49.46 61.86 63.95 66.47
Top 2 97.55 68.06 80.97 82.26 82.84
Top 5 97.55 86.37 95.03 94.23 94.88
Table 3: BERT & Llama2-chat: Accuracy of ConceptMapper in mapping a representation to the correct latent concept. See Table 1112 and 23 in the Appendix for detailed results.
4.3.2 PredictionAttributor
We question whether the salient input representation correctly represents the latent space of the output. This specifically evaluates PredictionAttributor. We calculate the number of times the representation of the most salient word/[CLS] token maps to the latent concept of the identical label as that of the prediction. We expect a high alignment at the top layers for PredictionAttributor to be correct. We do not include ConceptMapper when evaluating this and conduct the experiment using the training data only where we already know the alignment of a salient representation and the latent concept. Table 2 shows the results from the last four layers of 12-layered models, as well as the latter four layers of the Llama2-chat model (See Appendix Tables 910, and 22 for detailed results). For POS, we observed a successful match of above 90% for all models. We observed the mismatched cases and found them to be also of a compositional nature i.e. latent concepts comprised of semantically related words (see Appendix:Figure 9 for examples).
For sentiment, more than 99% of the time, the last layer’s salient representation maps to the predicted class label in 12-layered models, confirming the correctness of PredictionAttributor. The performance drop for the lower layer is due to the absence of class-based latent concepts in the lower layers i.e. concepts that comprised more than 90% of the tokens belonging to sentences of one of the classes. By comparison, the Llama2-chat model performs worse than the 12-layer model; however, it still achieves over 70% accuracy in its last layer.
4.3.3 ConceptMapper
We evaluate the correctness of ConceptMapper in mapping a test representation to the training data latent concepts. ConceptMapper trains using representations and their cluster IDs as labels. We randomly split this training data into 90% train and 10% test where the test data serves as the gold annotation of latent concepts. We train ConceptMapper using the train instances and measure the accuracy of the test instances. Table 3 presents the accuracy of POS using BERT, as well as for the Sentiment tasks using both BERT and Llama2-chat models (See Appendix Tables 111223). Observing Top-1 accuracy, the performance of ConceptMapper starts high (100%) for lower layers and drops to 84.19% and 68.24% for the last layer of BERT. A similar trend is observed for Llama2-chat, which exhibits 97.55% accuracy in the lower layer and decreases to 66.47% in the last layer. We found that the latent space becomes dense on the last layer. This is in line with Ethayarajh (2019) who showed that the representations of higher layers are highly anisotropic. This causes concepts to be close in the space. If true, the correct label should be among the top predictions of the mapper. We empirically tested it by considering the top two and top five predictions of the mapper, achieving a performance of up to 99.05% (BERT) for POS, 94.24% (BERT) and 94.88% (Llama2-chat) for Sentiment.
4.4 Faithfulness Evaluation
We consider a salient latent concept highlighted by LACOAT to be faithful to the prediction if the ablation of that latent concept causes a change in prediction performance. We define ablation of a latent concept as removing the information of that latent concept from the prediction vector i.e. [CLS]. We calculate the vector of a latent concept by averaging the training representations that belong to the latent concept. At inference time, we ablate the most salient latent concept for a prediction by subtracting the latent concept vector from the [CLS] representation of the last layer and then perform the prediction. The prediction change represents the percentage of predictions that were altered after the manipulation. As a baseline, we manipulate using a random vector of the same magnitude to that of the salient latent concept vector and report the change in prediction.
Table 4 reports the model’s accuracy and the percentage of altered predictions across all metrics. The results show that manipulating the [CLS] token representation using the LACOAT vector leads to significant performance drops and prediction changes across all datasets. In contrast, random vector manipulations have minimal impact on the model’s performance and predictions. These findings suggest that the LACOAT vector plays a crucial role in the model’s decision-making process. Comparing the results across different datasets, MNLI shows a relatively smaller drop in accuracy when manipulating the salient latent concept vector. We suspect this is due to the nature of the MNLI task, which requires reasoning over multiple sentences, with relevant information spread across multiple latent concepts. Nevertheless, the difference in results between the original accuracy and random vector confirms our hypothesis about the faithfulness of latent concepts.
Faithfulness Metrics
Dataset Setting Accuracy % Label Flip
Sentiment Original 96.31 -
LACOAT 55.91 43.98
Random 96.09 0.14
Toxicity Original 91.55 -
LACOAT 51.78 46.44
Random 91.93 0.13
MNLI Original 87.69 -
LACOAT 82.08 8.83
Random 88.12 0.55
Table 4: Faithfulness evaluation using the RoBERTa model. Original refers to the model’s performance without any manipulation, LACOAT represents the performance after subtracting the most salient latent concept vector from the [CLS] vector, and Random is the average performance of the model after subtracting five random vectors from the [CLS] vector.
5 Related work
The explainability methods can be approached by local explanations and global explanations targeting post-hoc analysis or introducing interpretability in the architecture Madsen et al. (2023); Sundararajan et al. (2017a); Denil et al. (2014); Selvaraju et al. (2020); Kapishnikov et al. (2021); Zhao and Aletras (2023); Kim et al. (2018); Ghorbani et al. (2019); Jourdan et al. (2023); Zhao et al. (2023); Ribeiro et al. (2016); Rajagopal et al. (2021). Lyu et al. (2023) provides a survey of explainability methods in NLP. LACOAT is a local explanation method providing post-hoc explanations given an input instance.
Previous work attempted to explain and interpret NLP models using human-defined concepts Kim et al. (2018); Abraham et al. (2022) and concepts extracted from hidden representations Zhao et al. (2023); Ghorbani et al. (2019); Rajani et al. (2020); "Geva et al. (2022). Zhao et al. (2023); Kim et al. (2018) worked on the global explanation based on a surrogate model. We provide local explanations and we ensure the faithfulness of latent concepts by extracting them directly from the hidden representation without any supervised training. Rajani et al. (2020) used k-nearest neighbors of the training data to identify erroneous correlations and misclassified instances.  Dalvi et al. (2022); Sajjad et al. (2022) analyzed latent concepts in their ability to represent linguistic knowledge. Our ConceptDiscoverer module is motivated by them. However, we propose a method to explain a model’s prediction using latent concepts.
6 Conclusion
7 Limitations
We discussed the limitations of LACOAT as follows:
  • While hierarchical clustering is better than nearest neighbor in discovering latent concepts as established by Dalvi et al. (2022), it has computational limitations and it can not be easily extended to a corpus of say 1M tokens. However, the assumptions that are taken in the experimental setup e.g. considering the maximum 20 occurrences of a word (supported by Dalvi et al. (2022)) work well in practice in terms of limiting the number of tokens and covering all facets of a majority of the words. Moreover, the majority of the real-world tasks have limited task-specific data and LACOAT can effectively be applied in such cases. In a contemporary work, Hawasly et al. (2024) recently targeted this limitation by exploring alternative algorithms such as leaders and k-means, demonstrating that these are viable, cheaper alternatives to the expensive agglomerative hierarchical clustering. Future explorations are needed to evaluate the correctness of the resulting clusters. For instance, k-means assume clusters to be spherical and similar assumption is enforced with the fixed distance threshold used in leaders. This limits their usefulness to identify clusters with irregular boundaries.
  • The concept discoverer is also constrained by a very small dataset. We conducted experiments to discover latent concepts with varying data sizes. The results indicate that when the data size is below 1000 instances, the discovered latent concepts tend to be sparse and lack coherence to form a clear concept.
  • The computational cost of LACOAT is higher than that of the IG method alone. While ConceptDiscoverer runs once per dataset as a pre-processing step, minimizing its impact on runtime during model inference, it still adds additional demand. Moreover, ConceptMapper is a linear classifier with an inference time of $O(d*c)$, where ${d}$ is the number of features and $c$ is the number of clusters. All other processes, aside from the IG part, have a linear cost.
  • For tasks requiring reasoning over multiple sentences, we observe that sometimes the LACOAT explanation’s are not clearly indicative of the reason of a prediction which might be based on some syntactic and semantic similarity between multiple input sentences. A possible solution to this is to consider hierarchical relationship between latent concepts in contrast to considering a flat structure among latent concepts. The underlying setup of ConceptDiscoverer supports this. However, comparing hierarchical structures requires further investigation beyond the scope of current work which provides a strong evidence towards faithful and human friendly explanations using training data latent space.
  • The human friendly explanations generated by PlausiFyer are prone to errors due to the inherent limitations of generative models such as positional bias Zheng et al. (2023); Khan et al. (2024), verbosity bias Huang et al. (2024), sensitivity to prompt style which hinders reproducibility. For instance, in this work, we observed that if we provide the predicted label to PlausiFyer, it generates an explanation describing the relationship between input, latent concept and label irrespective of whether the label is correct or not. A possible solution is to use multiple LLMs Verga et al. (2024); Wang et al. (2023); Badshah and Sajjad (2024) and LLMs agents Chen et al. (2024); Tan et al. (2024) that reasons over diverse explanations and reach to a better precise explanation of a latent concept.
Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI) and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada.
Appendix A Datasets
@@
Appendix B Finetuning Performance
We tuned several transformers BERT-base-cased, RoBERTa and XLM-RoBERTa. We used standard splits for training, development and test data that we used to carry out our analysis. The splits to preprocess the data are available through git repository. [^4] https://github.com/nelson-liu/contextual-repr-analysis See Table 6 and Table 7 for statistics and classifier accuracy. We present the results of Toxicity and MNLI in Appendix F and G.
@@ Appendix B Finetuning Performance
Appendix C Qualitative Evaluation - More Examples
C.1 Example for the Evolution of Concepts
@@
C.2 Adversarial Example of the Sentence in Figure 3(b)
@@
C.3 Correct Predicted Label with Incorrect Gold Label
@@
C.4 Incorrect Prediction in POS tagging Task
@@
Appendix D Module Specific Evaluation
D.1 ConceptDiscoverer - Compositional Concept Examples
@@
D.2 ConceptDiscoverer - Number of Clusters For Each Polarity in the Sentiment Classification Task
@@
D.3 ConceptMapper - Accuracy of ConceptMapper for the Sentiment Classification and POS Tagging task
@@
Appendix E Human Evaluation
E.1 LACOAT Effectiveness
@@
E.2 Comparison with other Methods
For comparison with other methods, we ask four annotators to rank 100 samples where they see the original input, gold label, predicted label, and explanations by three methods: LACOAT, IG and COCKATIEL. We consider the optimal settings for each explanation method. IG explanations are shown for layer 0 based on salient input features, while in layer 12, IG identifies only the [CLS] token as the most salient feature, which provides an implausible explanation and may negatively impact its human evaluation. COCKATIEL explanations are restricted to layer 12. LACOAT explanations span three layers (0, 6, and 12). By presenting LACOAT’s layerwise explanations, we offer a more comprehensive view of how the prediction evolves through the network, enhancing human evaluation.
The annotators are asked to rank each method from 1 to 3 in terms of usefulness in understanding the reason for the prediction where 1 implies the method was very useful while 3 implies it was not useful. The annotation allows for the annotator to rank multiple methods with the same usefulness rating, e.g. for a particular sample, both LACOAT and COCKATIEL can have the rank 1. This setting is intentional since the output of explanation methods is not directly comparable to each other due to the difference in their design and the targeted form and granularity of explanation. Table 1 presents the results. The results suggested that LACOAT is preferred or equally preferred by all annotators. The average Cohen’s $\kappa$ further shows a "fair agreement" between annotators and the consolidated ranking where consolidated ranking is the average rank across users.
Appendix F Toxicity Classification Task
F.1 Experimental Setup